Errata for

QUATERNIONIC SUBMANIFOLDS IN QUATERNIONIC

SYMMETRIC SPACES

(Tôhoku Math. Journ. 38 (1986), 513–538)

Hiroyuki Tasaki

October 16, 1999

p.518 Line 1: “⟨ι(X)ι, (Y)⟩′” should read “⟨ι(X), ι(Y)⟩′”